20160322 » History » Revision 16

Revision 15 (Boyd, Steven, 23 March 2016 08:59) → Revision 16/18 (Boyd, Steven, 23 March 2016 08:59)

h1. Ops/Comm Scheduling meeting 

 Location : CR11 R3 @ 15:30 pm on Tuesday 22nd 

 Phone access :    ReadyTalk number  
 US 866-740-1260 
 UK 0800-496-0576 
 Swiss 0800-70-53-51 
 Italy 800182592 
 Access Code: 2155402 

 h2. Agenda 

 * *Actions from last meeting* 

 * MOM report - Data-taking this week 
 ** Lessons learnt during the run 
 *** Communication to shift teams during interventions 
 ** Reminder on software updates during running 
 ** Summary of runs (MOM) attachment:March22MomReport.pdf 
 ** Data plots from recent shift 
 *** Detector plots: attachment:OnlinePIDplots.pdf  
 *** TOF (Durga)  
 *** EMR (Francois) 
 *** Tracker (Melissa) attachment:SPSummary.png 
 *** Cerenkov (?)  
 *** KL (?)  
 * DC report - activities in the Hall this week 
 ** Pressure control valve on SSU 
 ** SSD Compressor failure 
 ** Decay solenoid fridge instability : [attachment:Plot_of_Stability_of_Decay_Solenoid.pdf] 
 * Run plan for the next week 
 ** Progress on LiH runs 
 ** Shutdown for Easter/intercycle break 
 ** Work in the shutdown 
 * Issues to address after the run finished 
 ** Do we need to keep He running in the trackers during the shutdown? 
 ** Run control setting proton absorber 
 ** Run control goes straight to "Run Completed"    - typically on first run of the day 
 * AOB 

 * Chair of meetings for next week is DC/MOM 

 * DONM : 29/3/2016 

 h2. Minutes 

 h3. Present : Ryan, Paul, Colin, Sophie, Ken, Paul K, Henry, Pierrick, Craig, Steve, 
 Josef, Chris, Mike, John, Mark, Victoria, Joe, John N, Scott 

 *Actions from last meeting 

 ** Paul H identified the issue of recording pressure and material in the absorber volume in the CDB. Pierrick 
 says that we still waiting for Janusz to make the relevant table. *Action stands* 

 * MOM report 
 ** The MOM presented a summary of data-taking. It was shown that uptime during periods of data-taking (and minus issues 
 we could not affect such as ISIS down periods) is around 92%. Shifters were congratulated for this impressive performance. 
 ** Lessons learnt 
 *** Response of experts when shifters identify a problem was praised. It was recognised that communication between experts and shfiters could be a bit sharper in order to facilitate efficient data-taking. 
 *** Henry brought up the confusion between "standard" and "special" run tags and asked if there was a difference in settings. It was thought not, but the possibility of confusion was recognised. *Action Item :* the MOM should be clear in the run plan about what tag to use for which setting. 
 ** Plots. Ken reminded those responsible for detectors that they need to look at the data and check that their detector 
 is operating properly. They also need to produce plots showing this for each meeting. 
 ** Steve reminded everyone to go through the correct procedure when altering online software. This procedure involves filling out a software change control form (available on the elog) and getting permission from the MOM before doing anything. 
 ** John said that the physics devil isn't working with the new MAUS version. Scott said that he is looking into this now. 
 * DC report 
 ** A problem was identified with a pressure control valve that had not been installed correctly and required some hours to 
 fix. There was much discussion about this. The summary is that system owners should monitor work on their systems and be sure 
 that all work has been done, as there is a tendency for overloaded workers in the Hall to be given a blizzard of jobs and, 
 perhaps unsurprisingly, start work on one, then another and forget the first. Better oversight of jobs is needed. 
 ** A compressor on SSD developed a leak. John thinks it has something to do with the fact that we are till using demineralised 
 water and emphasised the need to change to non-demin supply (going back to a discussion of some months ago). He said he would 
 like to transfer to the ISIS cooling circuit, but they are reviewing the request as they have some constraints on their side. 
 Any intervention is likely to stop our cooling loop for up to half a day, which is problematic for the magnets. Colin pointed 
 out that the MIPO had decided not to do this water work. However if the use of de-min water is now damaging the compressors then 
 the situation has elevated and we should revisit the argument. Josef said that we don't know if the corrosion in the compressor 
 was the problem and won't until he can take the damaged compressor apart and investigate the issue. Action was left at this 
 point, with Josef to take the compressor apart and report the findings. 
 ** Mark reported on an intervention he made to the decay solenoid fridge settings to fix an oscillation in the pressure. This worked but is not ideal. Mike wants to get into the DSA at the next shutdown and go through the system carefully to see what 
 the problem actually was, as well as talk to the Linde engineer. 
 * Run plan for next week 
 ** LiH running continues. Hopefully we may just have the time to take all the data we need. 
 ** We plan to keep the Hall closed during the Easter break to stop people getting in. We will not be in controlled entry 
 though and the outer airlock door will just be locked with a key that can be accessed by ISIS personnel as well. *Action Item :* Steve should inform Alan Stevens about our plans. 
 ** The Hall plan which Colin has made should be updated with new actions (work in DSA by Mike, possible water work as well). 
 * Issues to address after the run 
 ** We decided that we did *not* need to keep He flowing in the trackers during the shutdown. 
 ** Access to proton absorber settings in the CDB is now available, so fixing proton absorber values by hand in Run Control is no longer necessary 
 ** Proton absorber itself seems to stick. *Action Item :* Henry to look into this 
 ** Run Control is still going straight to "Completed" on the first run of the day. Something to do with stale run settings in DATE which also seems to crash at this point. 
 ** Mike pointed out the discrepancy on the DS of the readback currents, set currents and currents displayed by EPICS. He would like Daresbury to have a look at the DS and calibrate so that we understand which of these currents is the right one. It was 
 stated that the currents displayed on EPICS are identical to those actually running in the DS, but we need to check.  
 *Action Item :* Henry to look into this 
 * AOB 
 ** Henry asked that we take all relevant information off the whiteboard in the MLCR and store it somewhere useful. THe whiteboard 
 keeps getting cleaned whenever anyone brushes against it....